OpenStudy (anonymous):

Use the process of elimination to solve this sytem. -x + 2y = 8 -x + 10y = 48 I just do not quite understand how to use the process of elimination

OpenStudy (anonymous):

2y=-x+8 y=-x/2+4

OpenStudy (anonymous):

is that the whole answer or is that how I start it?